OpenStudy (anonymous):

How does sin x cos pi/2 + cos x sin pi/2 simplify to cos x?

Vocaloid (vocaloid):

cos(pi/2) = 0 sin(pi/2) = 1 sin(x)*cos(pi/2) + cos(x)*sin(pi/2) = sin(x)*0 + cos(x)*1 = 0 + cos(x) = cos(x)

OpenStudy (anonymous):

THANK YOU